OpenStudy (anonymous):

Consider the functions f(x) = 2x + 1 and g(x) = x2 − 10. What is the value of (f − g)(5)? a.−4 b. 4 c.−11 d.26

myininaya (myininaya):

f(5)-g(5)=?

myininaya (myininaya):

f(5)=2(5)+1=10+1=11 g(5)=5^2-10=25-10=15 so f(5)-g(5)=11-15

OpenStudy (anonymous):

are you sleeping myin ?

OpenStudy (anonymous):

so it -11 no im not

myininaya (myininaya):

11-15 =-4 not -11

OpenStudy (anonymous):

oooo i got it i put the wrong number thanks